Solve.
What expression raised to the fourth power is 81x to the 12th power y to the 8th power z to the 16th power

Answer by SHUgrad05(58)

To solve this take the 4th root of (81x^12y^8z^16):
Note: When taking and exponent to a power, you multiply the exponents and when multiplying, you add exponents.
The 4th root of 81=3 (3*3*3*3 = 81)
The 4th root of x^12=x^3 (x^3*x^3*x^3*x^3 = x^12)
The 4th root of y^8=y^2 (y^2*y^2*y^2*y^2 = y^8)
The 4th root of z^16=z^4 (z^4*z^4*z^4*z^4 = z^16)
3x^3y^2z^4